sum of two consecutive odd integers is 144. What is the smaller odd integer? What is the larger odd integer?
Answer by cher781(8)

Smaller=x Larger=x+2
x+x+2=144
2x+2=144
2x+2-2=144-2
2x=142
2x/2=142/2
x=71
Smaller=71
Larger=73
Check: 71+73=144